# Visual Localization in 3D Maps: Comparing Point Cloud, Mesh, and NeRF Representations

### v1

@@ V-C Localization in Directions Unseen during Mapping
The initial recording was used to reconstruct a mapping database with only the front-facing camera, as shown in Fig. 9(a). The map was generated by traversing from point 1 to point 2, where 98 images were captured (illustrated in blue). We then attempted to localize to the map during a reverse traversal from points 2 to 1, with the results depicted using orange-colored camera views. Of the 87 query images, only 4 were successfully localized (when turning around the corner) due to minimal overlap in the image views. A similar pattern was observed with COLMAP; although the localization module was given a perfect reconstruction, it failed to localize any query images. HLoc performed slightly better than COLMAP due to its learning-based feature matching approach.
NeRF also demonstrated suboptimal performance in generating images from the reverse perspective, likely due to insufficient training views. In contrast, the mesh-base pipeline demonstrated good performance. It could effectively generate rendered images facing in the un-mapped direction (albeit with imperfections resulting in holes in the rendered images). Despite these visual limitations, mesh-rendered images achieved a localization rate of approximately .
@@
V-D Ablation Study: Changes in Environment
@@
In both locations, lighting variation between the 2-3 month period was clearly observed. The detailed results of localization performance are presented in Tab. IV. All three map representations show a small drop in retrieval and location rates. However, they can still localize around of query images, demonstrating the robustness of matching against one unified synthetic image database, aided by learning-based visual feature detector SuperPoint and matcher SuperGlue.
V-E Ablation Study: Feature Detection and Matching
@@
Tab. V presents quantitative results for this test, comparing several learning-based and classic feature detection and matching techniques. Among the learning-based feature detectors, SuperPoint, R2D2, and D2Net show promising performance by matching many features. However, traditional methods like SIFT and Akaze match fewer features, many of which are incorrect. Note that R2D2 produces descriptors of size 128, and D2Net generates descriptors of size 512. Given that the SuperGlue matcher is pre-trained with descriptors of size 256, we use the Nearest Neighbour (NN) matcher with a ratio threshold test for R2D2 and D2Net. SuperPoint and R2D2 are capped at 1024 features, while D2Net, a dense image feature extractor, naturally yields more feature points.
@@
VI Conclusions and Future Work
VI-A Conclusions
In this study, we introduced a cross-modal localization system tailored to the task of localizing a single camera image within a color 3D map. We demonstrated a pipeline to construct 3D prior maps using three distinct representations: point clouds, meshes, and NeRF. Each representation could then be used to synthesize RGB and depth images. Utilizing the color 3D map as input, our visual localization pipeline estimates the camera pose of a query image through a retrieval and matching process, leveraging learning-based descriptors and feature detectors. We conducted a comprehensive analysis of localization performance across these representations and discussed their respective merits. Additionally, we offered a benchmark comparison with two purely vision-based localization systems to situate our results within the wider field of visual localization. Notably, both point cloud and mesh representations achieve a localization accuracy of for query images, while the NeRF representation surpasses these, achieving a localization rate of .
VI-B Limitation and Future Work
VI-B1 Scene Change
@@
VI-B2 Synthetic Image
All three map representations encounter specific challenges in accurately rendering scenes with minimal texture details. This difficulty is particularly notable in 3D mapping and reconstruction, where effectively capturing the RGB texture details of ground and vegetation proves challenging. Furthermore, during the online localization of our system, images that predominantly contain low-texture objects often yield insufficient features for matching, resulting in inaccurate pose estimations. Addressing these challenges remains a significant area for improvement.

### v2

@@ 5.4 Localization in Directions Unseen during Mapping
The initial recording was used to reconstruct a mapping database with only the front-facing camera, as shown in Fig. 9(a). The map was generated by traversing from point 1 to point 2, where 98 images were captured (illustrated in blue). We then attempted to localize to the map during a reverse traversal from points 2 to 1, with the results depicted using orange-colored camera views. Of the 87 query images, only 4 were successfully localized (when turning around the corner) due to minimal overlap in the image views. A similar pattern was observed with COLMAP where it failed to localize any query images even though the localization module had an accurate SfM reconstruction. HLoc performed marginally better than COLMAP due to its learning-based feature matching approach.
NeRF generated blurrier images from reverse perspectives when trained exclusively with front-facing camera images, resulting in suboptimal localization performance. This issue is likely caused by insufficient training views when the path involves walking in a straight line. In contrast, the mesh-based pipeline demonstrated strong performance in this scenario. Mesh could effectively generate rendered images facing in the un-mapped direction (albeit with imperfections resulting in holes in the rendered images). Despite these visual limitations, mesh-rendered images achieved a localization rate of approximately .
@@
5.5 Ablation Study: Changes in Environment
@@
In both locations, lighting variation between the 2-3 month period was clearly observed. The detailed results of localization performance are presented in Tab. 4. All three map representations show a small drop in retrieval and location rates. However, they can still localize around of query images, demonstrating the robustness of matching against one unified synthetic image database, aided by learning-based visual feature detector SuperPoint and matcher SuperGlue.
5.6 Ablation Study: Feature Detection and Matching
@@
Tab. 5 presents quantitative results for this test, comparing several learning-based and classic feature detection and matching techniques. Among the learning-based feature detectors, SuperPoint, R2D2, and D2Net show promising performance by matching many features. However, traditional methods like SIFT and Akaze match fewer features, many of which are incorrect. Note that R2D2 produces descriptors of size 128, and D2Net generates descriptors of size 512. Given that the SuperGlue matcher is pre-trained with descriptors of size 256, we use the Nearest Neighbour (NN) matcher with a ratio threshold test for R2D2 and D2Net. SuperPoint and R2D2 are capped at 1024 features, while D2Net, a dense image feature extractor, naturally yields more feature points.
@@
6 Conclusions and Future Work
6.1 Conclusions
Using readily available 3D color maps generated from SLAM missions or TLS scanners, we can repurpose these maps for localization tasks. In this paper, we introduce a localization system specifically designed to localize a single camera image within 3D color maps by sampling and generating a synthetic image database. We first demonstrated a pipeline to construct 3D prior maps using three distinct representations: point clouds, meshes, and NeRF. Each representation is then used to synthesize RGB and depth images. Following this, we proposed a strategy to define the set of rendering poses that optimize the creation of a representative 3D map while retaining a minimal set of database images. The visual localization pipeline can estimate the camera pose of a query image through a retrieval and matching process, leveraging learning-based descriptors and feature detectors. We conducted a comprehensive analysis of localization performance across these representations and discussed their respective merits. Additionally, we offered a benchmark comparison with two purely vision-based localization systems to situate our results within the wider field of visual localization. Notably, both point cloud and mesh representations achieve a localization accuracy of for query images, while the NeRF representation surpasses these, achieving a localization rate of .
6.2 Limitation and Future Work
6.2.1 Scene Change
@@
6.2.2 Synthetic Image
All three map representations encounter specific challenges in accurately rendering scenes with minimal texture details. This difficulty is particularly notable in 3D mapping and reconstruction, where effectively capturing the RGB texture details of ground and vegetation proves challenging. Furthermore, during the online localization of our system, images that predominantly contain low-texture objects often yield insufficient features for matching, resulting in inaccurate pose estimations. To address this, we could prune the database images during online localization to retain only the most useful images and synthesizing new ones on-the-fly as needed.